# Current Address: ]Kinlin School, Fanshawe College, 1001 Fanshawe College Blvd., London, Ontario, Canada

### v2

@@ VII.2 Numerical Examples
Applying the procedures of Sec. VII.1 to model $\mathcal{M}^{(3)}_{A}$ [Eq. (97)], which is real and asymmetric, yields the results shown in Fig. 1; two different values of the Brody parameter $q$ (as indicated) are considered. The following parameter values and variables pertinent to model $\mathcal{M}^{(3)}_{A}$ were employed in the calculations: $\sigma_{\text{e}}=\ln(2)<1$, $\{c_{1}=c_{2}=c_{3}=1,c_{4}=-1\}$, $U\sim\text{Rayleigh}(\sigma_{R}=2)$, and $V=|T|$, where $T\sim\text{Normal}(\mu_{T}=1,\sigma_{T}^{2}=9)$. Numerics and theory are clearly consistent.
@@ VIII.1.1 Comment M1
The power-transformed exponential random variables $\left\{Y^{{p_{1}}\over{q+1}},Y^{{p_{2}}\over{q+1}},\ldots,Y^{{p_{n}}\over{q+1}}:n\in\mathbb{N}\right\}$ appearing in the definition of model (6) are generally distinct variables but they are not independent since they are merely different powers (i.e., different functions) of the same random variable $Y$. The matrix elements (which are different linear combinations of these dependent variables) are therefore generally dependent; it is in this sense that model (6) is a “correlated” random-matrix model. A secondary but nonetheless fundamental correlation in the model is that the exponents of $Y$ are themselves necessarily related (due to condition (8)). The net effect of this secondary correlation is that the spacing distribution ends up depending on only the distribution of the random variable $Y^{1\over{q+1}}$. The distribution(s) of any other random variable(s) the matrix element(s) could involve (e.g., the variables $\{U,V,X\}$ in model $\mathcal{M}^{(3)}_{A}$ [Eq. (97)]) do not affect the spacing distribution.
@@ VIII.1.2 Comment M2
If all members of $\mathbb{P}\equiv\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}:n\in\mathbb{N}\right\}$ in model (6) are constants, then the matrix elements of any cropped or pruned sub-classes of models are (in general) fully correlated in the sense that the values of all matrix elements can be obtained from knowledge of the value of any one matrix element. This perfect correlation is generally destroyed when some members of $\mathbb{P}$ are random variables. For example, if $X=0$ in (trimmed) model (97), then it is clear that the values of the off-diagonal matrix elements cannot be obtained from knowledge of the values of the diagonal matrix elements. Sub-classes of models (derived from model (6)) that are not cropped or pruned are, in general, not perfectly correlated. The takeaway is: condition (8) does not necessarily imply that the matrix elements are fully correlated. This subtle but important point applies to all $2\times 2$ random-matrix models satisfying condition (8), whether they derive from model (6) or not (e.g., sub-class J (47)).
@@ VIII.1.3 Comment M3
do not derive from model (6) but nonetheless satisfy condition (8) and hence have Brody spacing distributions. In models (107), the random variable $X$ is independent of $Y$ and can have any arbitrary continuous distribution $f_{X}(x)$, and $c(q)$ is a real $q$-dependent constant that can be arbitrarily prescribed. The other model-specific quantities are as follows: in model (107c), $g(X)$ is any real function of $X$ having the property $g(X)\neq 0$ for all values of $X$; in model (107f), the real exponents $\{a,b\}$ satisfy the condition $a+b=1$ but are otherwise arbitrary; and in model (107i), the diagonal functions $\{d_{1}(X),d_{2}(X)\}$ are any two real functions of $X$ that satisfy $d_{1}(X)-d_{2}(X)=1$ ($\forall X$). In each of models (107c)-(107i), the off-diagonal matrix elements are independent of the diagonal matrix elements. Note that the off-diagonal elements of $\mathcal{M}_{D3}$ can be real or pure imaginary; the eigenvalues of $\mathcal{M}_{D3}$ are however always real.
It is worth mentioning that correlated “additive” models satisfying condition (8) can also be constructed (two simple examples are provided in Appendix D). Incidentally, the generalizations of Cases I and III of sub-class B given in Sec. IV.2.3 are also models that are not directly derivable from model (6), but the same comment as above applies to these two cases as well.
@@ VIII.1.4 Comment M4
$\mathcal{D}(\mathcal{M})=kV^{2},~{}~{}k\neq 0,~{}V\sim\text{Weibull}(\sigma=\sigma(\sigma_{1},\sigma_{2},\ldots,\sigma_{m}),\tau=q+1),$ (108)
where $\sigma=\sigma(\cdots)$ denotes the fact that the Weibull scale parameter $\sigma$ can depend on any distributional scale parameters the random-matrix model may involve.
@@
It can be shown (see Appendix E) that the eigenvalue spacing (c.f., Eq. (10))
$S=\left(U^{2}+V^{2}\right)^{1\over{q+1}}\sim\text{Weibull}\left(\sigma=2\sigma^{2}_{R},\tau=q+1\right),$ (113)
and that the distribution of $Z\equiv{S/\mu_{S}}$ (i.e., the mean-scaled spacing distribution) is again the Brody distribution. Comparing (12) and (113), it is then clear that model $\mathcal{M}_{R}$ satisfies condition (8) (with $k=1$) only in the special case when $\sigma_{R}=\sqrt{\sigma_{\text{e}}/2}$. Therefore, a key point is that the above example satisfies condition (108) (with $k=1$ and $V=R^{2\over{q+1}}$) but does not (in general) satisfy condition (8). The reason is that condition (8) [with $f_{Y}(y)$ as given by Eq. (7)] and condition (108) are not identical; the former is a special case of the latter.
@@ VIII.2.2 Comment P2
The $q=0$ limit of model (6) can (subject to condition (8)) generate a variety of new classes of $2\times 2$ (correlated) random-matrix models having Poisson spacing statistics. For example, the eigenvalue spacings of real asymmetric random matrices of the form (c.f., Sec. IV.2.2)
@@ VIII.2.4 Comment P4
Pruned and cropped models may seem unphysical due to the (conditionally) fully-correlated nature of the matrix elements (c.f., Comment M2). Random matrix models having highly-correlated entries are actually not a novelty in physics (see, for example, Refs. [74, 75, 76, 77] and references contained therein). The success of classical RMT in physics (and other fields) does not rely on all of its specific model assumptions. As mentioned in the Introduction, there is a large body of mathematical research that rigorously shows that the classical results of RMT do not depend on the specific nature of the matrix elements (including whether or not the entries are correlated) but rather only on the symmetry class of the matrix (e.g., real symmetric, complex Hermitian, etc.). This is the celebrated spectral universality that was originally conjectured by Wigner, Dyson, Gaudin, and Mehta [11]. The classical independence assumption used in the construction of the GOE is an assumption that simplifies the mathematical analysis and allows for analytical results but has no direct connection to quantum mechanics since the matrix elements of quantum Hamiltonian systems are typically not independent (see, for example, Refs. [78, 79]). In an inverse-problem setting where the end goal is to construct models having a specific spacing distribution, the matter of whether the matrix elements are uncorrelated, partially correlated, or fully correlated, is therefore a not a primary concern. Ideally, all three correlation classes should be considered in turn. It may seem backwards to start with fully-correlated models (instead of uncorrelated ones), but it turns out that (in the case of the Brody distribution) the fully-correlated case is the simplest starting point for model construction, both mathematically and conceptually.
It is worth re-iterating (c.f., the paragraph preceding Sec. IV.1) that the full correlation (conditionally) present in pruned and cropped models can (if desired) be selectively destroyed by employing random variables in the summand exponents (c.f., Comment M2) and/or making use of the functions $\{g_{1}(X),g_{2}(X),h(T)\}$ in the general model (6). For reference, Table 1 lists the models given in the current paper that possess independent matrix elements. In general, mutual independence between any pair of matrix elements can be effected by judicious placement of random functions. For instance, it is clear from inspection of Table 1 that no example was given that possessed mutually independent off-diagonal elements. This particular independence structure can be easily effected in many ways using random functions. A simple example is:
$\mathcal{M}_{L}=\left(\begin{array}[]{cc}Y^{1\over{q+1}}&{1\over 2}\sin^{2}(2qV)\\ {1\over 2}Y^{2\over{q+1}}&~{}2\sin^{2}(qV)Y^{1\over{q+1}}\end{array}\right),$ (122)
@@ A.1 The Weibull Distribution
Let the random variable $W\sim\text{Weibull}(\sigma,\tau)$, that is, let $W$ have a Weibull distribution with scale parameter $\sigma>0$ and shape parameter $\tau>0$. The p.d.f. of $W$ is then given by:
$f_{W}(w;\sigma,\tau)={\tau\over\sigma}w^{\tau-1}\exp\left(-{w^{\tau}\over\sigma}\right),~{}w\geq 0.$ (123)
@@
$\mu_{W}(\sigma,\tau)=\left({1\over\sigma}\right)^{-{1/\tau}}\Gamma\left(1+{1\over\tau}\right).$ (124)
@@ A.2 Generalized Gamma Distribution
$V^{{}^{p}}\sim\text{GG}\left(\ell^{{}^{p}},{\omega\over p},{\Omega\over p}\right).$ (128)
@@ A.3 Gamma Distribution
$X^{{}^{p}}\sim\text{GG}\left(\ell=(\sigma_{\text{g}})^{{}^{p}},\omega=\xi/p,\Omega=1/p\right).$ (130)
Appendix B Discriminant Constants
For reference, the discriminant constants for sub-classes A and B are as follows:
$k_{A}=(c_{1}-c_{4})^{2}+4c_{2}c_{3},$ (131a)
$k^{(j)}_{B}=\mathcal{C}_{j},~{}j=\{1,2,3\},$ (131b)
[…]
Appendix C Derivation of Distribution (102)
The derivation of distribution (102) closely parallels the derivation of the Brody distribution given in Sec. III. The task in this appendix is to determine the distribution of the random spacing $S$ as given by Eq. (11) with $Y\sim\text{Gamma}(\sigma_{\text{g}},\xi=2)$ (instead of $Y\sim\text{Exp}(\sigma_{\text{e}})$). This can be efficiently accomplished by exploiting property (130), which immediately yields:
@@ Appendix C Derivation of Distribution (102)
Appendix D Additive-Type Correlated Models: Two Simple Examples
@@
Appendix E Derivation of Result (113)
@@
$f_{R}(r;\sigma_{R})={1\over\sigma^{2}_{R}}r\exp\left(-r^{2}/2\sigma^{2}_{R}\right),~{}r\geq 0.$ (164)

### v3

@@ VII.2 Numerical Examples
Applying the procedures of Sec. VII.1 to model $\mathcal{M}^{(3)}_{A}$ [Eq. (94)], which is real and asymmetric, yields the results shown in Fig. 1; two different values of the Brody parameter $q$ (as indicated) are considered. The following parameter values and variables pertinent to model $\mathcal{M}^{(3)}_{A}$ were employed in the calculations: $\sigma_{\text{e}}=\ln(2)<1$, $\{c_{1}=c_{2}=c_{3}=1,c_{4}=-1\}$, $U\sim\text{Rayleigh}(\sigma_{R}=2)$, and $V\sim\text{THN}(\mu_{\text{h}}=1,\sigma_{\text{h}}^{2}=9)$, that is, $V$ has a truncated half-normal distribution with location parameter $\mu_{\text{h}}=1$, scale parameter $\sigma_{\text{h}}=3$, and (truncated) support $v\in[1,\infty)$. Numerics and theory are clearly consistent.
@@ VIII.1.1 Comment M1
The power-transformed exponential random variables $\left\{Y^{{p_{1}}\over{q+1}},Y^{{p_{2}}\over{q+1}},\ldots,Y^{{p_{n}}\over{q+1}}:n\in\mathbb{N}\right\}$ appearing in the definition of model (6) are distinct variables but they are not independent since they are merely different powers (i.e., different functions) of the same random variable $Y$. The matrix elements (which involve different linear combinations of these dependent variables) are therefore generally dependent; it is in this sense that model (6) is a “correlated” random-matrix model. A secondary but nonetheless fundamental correlation in the model is that the exponents of $Y$ are themselves necessarily related (due to condition (8)). The net effect of this secondary correlation is that the spacing distribution ends up depending on only the distribution of the random variable $Y^{1\over{q+1}}$. The distribution(s) of any other random variable(s) the matrix element(s) could involve (e.g., the variables $\{U,V,X\}$ in model $\mathcal{M}^{(3)}_{A}$ [Eq. (94)]) do not affect the spacing distribution.
@@ VIII.1.2 Comment M2
If all members of $\mathbb{P}\equiv\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}:n\in\mathbb{N}\right\}$ and $\mathbb{K}\equiv\{k_{1},k_{2},\ldots,k_{4n+4}:n\in\mathbb{N}\}$ in model (6) are constants, then the matrix elements of any cropped or pruned sub-classes of models are (in general) fully correlated in the sense that the values of all matrix elements can be obtained from knowledge of the value of any one matrix element. This perfect correlation is generally destroyed when some members of $\mathbb{P}$ and/or $\mathbb{K}$ are random variables. For example, if $X=0$ in (trimmed) model (94), then it is clear that the values of the off-diagonal matrix elements cannot be obtained from knowledge of the values of the diagonal matrix elements. Likewise, the value of the lower off-diagonal matrix element in model (105) cannot be obtained from sole knowledge of the value of the upper off-diagonal matrix element. Sub-classes of models (derived from model (6)) that are not cropped or pruned are, in general, not perfectly correlated; this is true even when all members of $\mathbb{P}$ and $\mathbb{K}$ are constants. The takeaway is: condition (8) does not necessarily imply that the matrix elements are fully correlated. This subtle but important point applies to all $2\times 2$ random-matrix models satisfying condition (8), whether they derive from model (6) or not (e.g., sub-class J (44)).
@@ VIII.1.3 Comment M3
do not derive from model (6) but nonetheless satisfy condition (8) and hence have Brody spacing distributions. In models (112), the random variable $X$ is independent of $Y$ and can have any arbitrary continuous distribution $f_{X}(x)$, and $c(q)$ is a real $q$-dependent constant that can be arbitrarily prescribed. The other model-specific quantities are as follows: in model (112c), $g(X)$ is any real function of $X$ having the property $g(X)\neq 0$ for all supported values of $X$; in model (112f), the real exponents $\{a,b\}$ satisfy the condition $a+b=1$ but are otherwise arbitrary; and in model (112i), the diagonal functions $\{d_{1}(X),d_{2}(X)\}$ are any two real functions of $X$ that satisfy $d_{1}(X)-d_{2}(X)=1$. In each of models (112c)-(112i), the off-diagonal matrix elements are independent of the diagonal matrix elements. Note that the off-diagonal elements of $\mathcal{M}_{D3}$ can be real or pure imaginary; the eigenvalues of $\mathcal{M}_{D3}$ are however always real.
It is worth mentioning that correlated “additive” models satisfying condition (8) can also be constructed (two simple examples are provided in Appendix C). Incidentally, the generalizations of Cases I and III of sub-class B given in Sec. IV.2.3 are also models that are not directly derivable from model (6), but the same comment as above applies to these two cases as well.
@@ VIII.1.4 Comment M4
$\mathcal{D}(\mathcal{M})=kL^{2},~{}~{}k\neq 0,~{}L\sim\text{Weibull}(\theta=\sigma(\sigma_{1},\sigma_{2},\ldots,\sigma_{l}),\tau=q+1),$ (113)
where $\theta=\sigma(\cdots)$ denotes the fact that the Weibull scale parameter $\theta$ can depend on any distributional scale parameters $\{\sigma_{1},\sigma_{2},\ldots,\sigma_{l}:l\in\mathbb{N}\}$ the random-matrix model may involve.
@@
It can be shown (see Appendix D) that the eigenvalue spacing (c.f., Eq. (10))
$S=\left(U^{2}+V^{2}\right)^{1\over{q+1}}\sim\text{Weibull}\left(\theta=2\sigma^{2}_{R},\tau=q+1\right),$ (118)
and that the distribution of $Z\equiv{S/\mu_{S}}$ (i.e., the mean-scaled spacing distribution) is again the Brody distribution. Comparing (12) and (118), it is then clear that model $\mathcal{M}_{R}$ satisfies condition (8) (with $k=1$) only in the special case when $\sigma_{R}=\sqrt{\sigma_{\text{e}}/2}$. Therefore, a key point is that the above example satisfies condition (113) (with $k=1$ and $L=R^{2\over{q+1}}$) but does not (in general) satisfy condition (8). The reason is that condition (8) [with $f_{Y}(y)$ as given by Eq. (7)] and condition (113) are not identical; the former is a special case of the latter.
@@ VIII.2.2 Comment P2
The $q=0$ limit of model (6) can (subject to condition (8)) generate a variety of new classes of $2\times 2$ (correlated) random-matrix models having Poisson spacing statistics. For example, the eigenvalue spacings of real asymmetric random matrices of the (pruned) form (c.f., Sec. IV.2.2)
@@
Although not of primary interest in the current paper, it should be noted that the $q=0$ limit of model (6) with $f_{Y}(y)$ given by (106) (rather than (7)) can generate novel classes of real or complex $2\times 2$ random matrices possessing real or complex eigenvalues with semi-Poissonian spacings (hermiticity is again not a requirement when the matrices are complex).
@@ VIII.2.3 Comment P3
Analogous comments apply to the $q=1$ limit of model (6) with $f_{Y}(y)$ given by (106) (rather than (7)) which can generate novel classes of real or complex $2\times 2$ random matrices possessing real or complex eigenvalues whose spacings follow the Ginibre distribution (109). This is interesting, both mathematically and physically, since it clearly shows that the elements of a $2\times 2$ random matrix need not be IID and/or Gaussian-distributed nor does the matrix need to be complex in order for its eigenvalues (which also need not be complex) to have a Ginibre spacing distribution (and hence cubic level repulsion). Moreover, the presence or absence of hermiticity/symmetry is irrelevant (at least when correlations between matrix elements are permitted). The preceding are fundamentally important points that (to the author’s knowledge) have not been discussed in the literature.
@@ VIII.2.4 Comment P4
Pruned and cropped models may seem unphysical due to the (conditionally) fully-correlated nature of the matrix elements (c.f., Comment M2). Random matrix models having highly-correlated entries are actually not a novelty in physics (see, for example, Refs. [74, 75, 76, 77, 78] and references contained therein). The success of classical RMT (in physics and other fields) does not rely on all of its specific model assumptions. As mentioned in the Introduction, there is a large body of mathematical research that rigorously shows that the classical results of RMT do not depend on the specific nature of the matrix elements (including whether or not the entries are correlated) but rather only on the symmetry class of the matrix (e.g., real symmetric, complex Hermitian, etc.). This is the celebrated spectral universality that was originally conjectured by Wigner, Dyson, Gaudin, and Mehta [11]. The classical independence assumption used in the construction of the GOE is an assumption that simplifies the mathematical analysis and allows for analytical results but has no direct connection to quantum mechanics since the matrix elements of quantum Hamiltonian systems are typically not independent (see, for example, Refs. [79, 80]). The matter of whether the matrix elements are uncorrelated, partially correlated, or fully correlated, is therefore not a general concern.
It is worth re-iterating that the full correlation (conditionally) present in pruned and cropped models can (if desired) be selectively destroyed by employing random variables in the summand exponents and/or coefficients (c.f., Comment M2) and/or making use of the functions $\{g_{1}(X),g_{2}(X),h(T)\}$ in the general model (6) (c.f., the paragraph preceding Sec. IV.1). For reference, Table 1 lists the models given in the current paper that possess independent matrix elements. In general, mutual independence between any pair of matrix elements can be effected by judicious placement of random functions. For instance, it is clear from inspection of Table 1 that no example was given that possessed mutually independent off-diagonal elements. This particular independence structure can be easily effected in many ways using random functions. A simple example is:
$\left(\begin{array}[]{cc}Y^{1\over{q+1}}&{1\over 2}\sin^{2}(2qV)\\ {1\over 2}Y^{2\over{q+1}}&~{}2\sin^{2}(qV)Y^{1\over{q+1}}\end{array}\right),$ (127)
@@ A.1 The Weibull Distribution
Let the random variable $W\sim\text{Weibull}(\theta,\tau)$, that is, let $W$ have a Weibull distribution with scale parameter $\theta>0$ and shape parameter $\tau>0$. The p.d.f. of $W$ is then given by:
$f_{W}(w;\theta,\tau)={\tau\over\theta}w^{\tau-1}\exp\left(-{w^{\tau}\over\theta}\right),~{}w>0.$ (128)
@@
$\mu_{W}(\theta,\tau)=\left({1\over\theta}\right)^{-{1/\tau}}\Gamma\left(1+{1\over\tau}\right).$ (129)
@@ A.2 Generalized Gamma Distribution
$V^{p}\sim\text{GG}\left(\ell^{p},{\omega\over p},{\Omega\over p}\right).$ (133)
@@ A.3 Gamma Distribution
$X^{p}\sim\text{GG}\left(\ell=(\sigma_{\text{g}})^{p},\omega=\xi/p,\Omega=1/p\right).$ (135)
Appendix B Derivation of Distribution (107)
The derivation of distribution (107) closely parallels the derivation of the Brody distribution given in Sec. III. The task in this appendix is to determine the distribution of the random spacing $S$ as given by Eq. (11) when $Y\sim\text{Gamma}(\sigma_{\text{g}},\xi=2)$ (instead of $Y\sim\text{Exp}(\sigma_{\text{e}})$). This can be efficiently accomplished by exploiting property (135), which immediately yields:
@@ Appendix B Derivation of Distribution (107)
Appendix C Additive-Type Correlated Models: Two Simple Examples
@@
Appendix D Derivation of Result (118)
@@
$f_{R}(r;\sigma_{R})={1\over\sigma^{2}_{R}}r\exp\left(-r^{2}/2\sigma^{2}_{R}\right),~{}r>0.$ (168)